9x+3y=27
12x-3y=36

Answer by checkley77(12844)

9x+3y=27
12x-3y=36 add these 2 equations.
-------------------
21x=63
x=63/21
x=3 ans.
9*3+3y=27
27+3y=27
3y=27-27
3y=0
y=0 ans
Proof:
12*3-3*0=36
36=36